OpenStudy (anonymous):

How would I factor this??? 8c2 - 26c + 15

OpenStudy (anonymous):

8c^2 - 26c + 15

OpenStudy (mathstudent55):

1. Try to factor out a common factor out of all terms if there is one.

OpenStudy (anonymous):

I don't think there is one.

OpenStudy (anonymous):

At least not one small enough to use in this equation.

OpenStudy (mathstudent55):

Correct. 2. This is a trinomial of the form ax^2 + bx + c Multiply a and c together. In this case it's 8 * 15

OpenStudy (mathstudent55):

ac = 8 * 15 =

OpenStudy (anonymous):

120

OpenStudy (mathstudent55):

Good. 3. Come up with two factors of ac that add to b. In this case you need two factors of 120 that add up to -26 Since 120 is positive and -26 is negative, the numbers you need are both negative.

OpenStudy (anonymous):

So I need negative factors of 120?

OpenStudy (mathstudent55):

Yes. You need two factors of 120; they must both be negative; they need to multiply to 120 and add to -26.

OpenStudy (anonymous):

-4 and -30!

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

-6 and -20

OpenStudy (mathstudent55):

Right

OpenStudy (mathstudent55):

4. Call these factors p and q. Now rewrite the trinomial with the middle term (bx) broken up into two terms (px + qx), ax^2 + px + qx + c In this case you get 8c^2 - 20c - 6c + 15 If you wrote 8c^2 - 6c - 20c + 15, it also works.

OpenStudy (mathstudent55):

5. Now factor by parts. Factor a common factor out of the first two terms, and factor a common factor out of the last two terms. 8c^2 - 20c - 6c + 15 =4c(2c - 5) - 3(2c - 5)

OpenStudy (mathstudent55):

6. Factor out the common term. 4c(2c - 5) - 3(2c - 5) Here, the common term is (2c - 5), so we factor it out: (2c - 5)(4c - 3) Now it's factored.

OpenStudy (mathstudent55):

This is called factoring by parts.

OpenStudy (anonymous):

for part 5, where did (2c-5) come from?

OpenStudy (anonymous):

Ohhh I think I know. 2 comes from 8 (because 4*2 = 8) and the c come from the 8. Then the 5 comes from the 20. Right?

OpenStudy (mathstudent55):

Right. Then there is another 2c - 5 that comes from -6c + 15

OpenStudy (anonymous):

Ohh. Makes sense.

OpenStudy (anonymous):

Now I get it. I guess I just need to get used to this. I haven't fully caught on to factoring stuff yet.

OpenStudy (anonymous):

Thank you for all your patience and help!!!

OpenStudy (mathstudent55):

This method of factoring is good for a trinomial like yours which has a number other than 1 with the x^2 term. In your case you had 8x^2. If you have a simpler trinomial which starts with just x^2, there is a simpler way of doing it.

OpenStudy (mathstudent55):

You're welcome.